OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 . (1 point) y - 6 = 1/2 (x + 3) y = 1/2 x - 6 y + 3 = 1/2 (x - 6) x - 2y = 12

OpenStudy (anonymous):

@pooja195 Can you help me please??

OpenStudy (anonymous):

Slope of 1/2 is the number in front of the x after you convert the form to y=mx+b

OpenStudy (anonymous):

So how would you convert the first one?

OpenStudy (kkutie7):

y=m(x-x1)+y1 m equals the slope. x1= x-coordinate and y1=y-coordinate

OpenStudy (kkutie7):

the equation can be written as y-y1=m(x-x1) too

OpenStudy (anonymous):

because all the equations are in different forms, you kinda just have to pick one of these forms and go with it.

OpenStudy (kkutie7):

@battlealex88 plug the numbers into the equation and it will be clear to you. y1=-3, x1=6, and m=1/2 y-y1=m(x-x1)